# Exercising the CCPA Opt-out Right on Android: Legally Mandated but Practically Challenging

### v5

@@ 3.2. Methods for Opting Out
The CCPA Regulations outline several methods through which businesses must allow consumers to exercise their opt-out right. [^11] CCPA Regulations, §7026(a). As a general guideline, the CCPA Regulations specify that businesses shall consider the methods by which they interact with consumers, the manner in which they collect the personal information that they make available to third parties, available technology, and ease of use by the consumer when determining which methods consumers may use to submit requests to opt out of selling and sharing with at least one method reflecting the manner in which the business primarily interacts with the consumer.
@@
For websites the CCPA specifies that businesses shall provide a “clear and conspicuous link on the business’ internet homepages, titled ‘Do Not Sell or Share My Personal Information,’ to an internet web page that enables a consumer, or a person authorized by the consumer, to opt out of the sale or sharing of the consumer’s personal information.” [^12] CCPA, §1798.135(a)(1) (emphasis added). In this regard, cookie banners or cookie controls, are not by themselves sufficient for enabling consumers to opt out from selling and sharing of personal information because those practices are not identical to cookie practices. [^13] CCPA Regulations, §7026(a)(4). While there may be some overlap—for example, disabling targeting cookies may also reduce selling and sharing of personal information—the CCPA requires a dedicated method for submitting requests to opt out of the selling and sharing of personal information.
@@
In addition to links the CCPA Regulations mandate that businesses respect opt-out preference signals: “A business that collects personal information from consumers online shall, at a minimum, allow consumers to submit requests to opt-out of sale/sharing through an opt-out preference signal and at least one of the following methods: an interactive form accessible via the ‘Do Not Sell or Share My Personal Information’ link, the Alternative Opt-out Link, or the business’s privacy policy if the business processes an opt-out preference signal in a frictionless manner.” [^14] CCPA Regulations, §7026(a)(1). The Alternative Opt-out Link means a link combining the two separate “Do Not Sell or Share My Personal Information” and “Limit the Use of My Sensitive Personal Information” links into one “Your Privacy Choices,” or, “Your California Privacy Choices” link per CCPA Regulations, §7001(b) and §7015. The idea is that businesses should (1) prominently notify consumers of the selling and sharing of personal information and how they can opt out and (2) process their opt-out preferences via automated opt-out preference signals. If a business processes opt-out preference signals in a frictionless manner, it is not required to provide any opt-out links. [^15] CCPA Regulations, §7025(e). Businesses are required to treat opt-out preference signals as a valid request to opt out of the sale or sharing of personal information for a browser or device and any consumer profile associated with that browser or device, including pseudonymous profiles. [^16] CCPA Regulations, §7025(c)(1).
@@ 3.3. Opting Out via GPC
The most important privacy preference signal for purposes of the CCPA is GPC, which is a draft standard at the W3C (Zimmeck et al., 2025). GPC signals are transmitted from users’ clients to servers via request headers. Servers can also query a user agent’s GPC status via the GPC DOM property.
GPC can be set to 1 (header) or true (DOM property) to express a user’s opt-out as defined by applicable law in the jurisdiction of the user. Thus, GPC can be used in multiple jurisdiction and will have the meaning that a particular law or regulation attaches to it. GPC aims to enable users to exercise their opt-out rights beyond individual websites and apps for sets of sites or apps.
Recent studies evaluating the usability of GPC suggest that most users understand what it does and that they would turn it on if it were more broadly available (Zimmeck et al., 2023, 2024b). Its rate of adoption by website operators appears to increase slowly and steadily estimated at a few percentage points per year (Hausladen et al., 2025). Compliance with GPC is required in California (since January 2021) (Becerra, 2021), Colorado (since July 2024) (Colorado Department of Law, 2024), Connecticut (Office of the Attorney General, 2025) (since January 2025), and New Jersey (New Jersey Division of Consumer Affairs, 2025) (since July 2025). The CCPA’s opt-out right and GPC apply not only to websites but also to mobile apps.
@@ 3.4. Opting Out via GPC on Mobile Apps
The Office of the California Attorney General brought various enforcement actions against mobile app operators for not respecting consumers’ opt-out right (State of California Department of Justice, 2022b, a, 2025). Consumers have a right to opt out on every “device,” which the CCPA defines as “any physical object that is capable of connecting to the internet, directly or indirectly, or to another device.” [^17] CCPA, §1798.140(o). Thus, the International Association of Privacy Professionals, a nonprofit and non-advocacy membership association of privacy professionals, recommends for mobile apps that “[a]t a minimum, mobile apps should […] include a functional ‘Do Not Sell or Share My Personal Information’ link in the app […]” (Desai and Zweifel-Keegan, 2023).
However, opting out via GPC on mobile devices presents a challenge. The core problem is the absence of a native and platform-level GPC setting on Android (and, for that matter, on any other platform beyond the web). This technical gap is in conflict with the CCPA Regulations, which mandate that “[t]hrough an opt-out preference signal, a consumer can opt-out of sale and sharing of their personal information with all businesses they interact with online without having to make individualized requests with each business.” [^18] CCPA Regulations, §7025(a) (emphasis added). This discrepancy between the law and the technical reality raises crucial questions: How are app operators expected to facilitate GPC opt-out preferences without platform-level support. Do they even anticipate receiving such signals? On a fundamental level, as mobile apps are using the web for their client–server communication, just as websites do, it is principally possible for them to process GPC signals in the same way as websites. Thus, our study investigates to which extent this is actually the case.
@@ 4.1. App Dataset
For our evaluation we downloaded a set of popular free apps from the US Google Play store in August 2023. Following previous work (Scoccia et al., 2020; Shuba and Markopoulou, 2020; Han et al., 2020), we identified the apps based on Google’s lists of popular apps in 49 app categories on the Google Play store. Each list contained 45 apps. The 49 app categories include 17 game categories (Google, 2024b). The sets of apps in the different categories are disjoint as each app is assigned exactly one category on the Google Play store (Google, 2024b). The 45 apps listed for each of the 49 categories are “heavily influenced by popularity” (Google, 2024a). [^19] Apps in the Communication and Productivity categories had the most installs while those in the Events category had the least. Appendix Figure 12 shows the app install ranges for the apps in all 49 app categories. We selected popular free apps for our study because those are most likely to rely on an ad-based business model that implies selling personal information or sharing such with third parties (Section 3.1). Using Raccoon (Onyxbits, 2022), we successfully downloaded 1,896/2,205 (86%) of the listed apps. The remaining apps were incompatible with our test device, removed from the Play Store before we could download them, or not successfully downloaded. We made two download attempts per app. For our GPC opt-out evaluation we successfully performed a dynamic analysis of 1,811/1,896 (96%) apps resulting in our app dataset. For the remaining apps we did not collect any or only little network traffic, for example, due to apps’ crashing during the analysis.
@@ 4.2. Measuring Opt-out Effects
Figure 2. Exercising the opt-out right via GPC. 1. An HTTP request from the user’s Android device is sent with a GPC header. By sending this header the user is exercising the opt-out right. 2. The HTTP response contains a script for setting a privacy flag client-side. By identifying such a privacy flag and its value we determine whether the opt-out has been respected. App operators can implement the logic for processing privacy flags themselves or use the library of a CMP. 3–4. A third-party ad network queries the privacy flag via its ad tag that the app operator integrated in its app. 5. The flag may be attached to ad requests from the user to the ad network, which is the second point that we use to determine whether our opt-out has been respected. The ad network can also act directly upon receiving GPC signals, which are attached to all requests as headers. 6. The ad network responds to ad requests from opted out users with an ad not based on sold or shared personal information.:
@@
We aim to measure the effects of opting out—i.e., whether an app stopped selling or sharing personal information—in three ways: we measure whether an app (1) sends to ad networks CCPA privacy flags that are set to opt-out (such as Vungle’s ccpa status set to opted out (Vungle, 2025)), makes fewer connections to ad tracking domains, and (3) discloses fewer device identifiers. As to the first measure, it is the responsibility of ad networks to provide functionality in their ad libraries for app operators to identify ad requests to which the CCPA applies. If an app operator enables such functionality, a privacy flag should be set on the user’s device and attached to outgoing ad requests from the app to the third party, as shown in Figure 2. Thus, as privacy flags are generally communicated client-to-server, for example, ad networks retrieve the US Privacy String from an Android’s device local SharedPreferences (IAB Tech Lab, 2020), we can use client-side analysis to determine our opt-out status. If a third party receives an ad request with a privacy flag, it knows to discard personal information upon receipt and serve an ad not based on sold or shared personal information.
@@
Beyond changing privacy flag values, a successful opt-out can manifest itself through other observable changes in an app’s behavior. An app can comply by not making any ad calls. Thus, monitoring an app’s connections to third-party ad tracking domains serves as another indicator for opt-out adherence. Further, an app might connect to third-party ad domains but stop the sharing of identifying data. Thus, we also examine the impact of opt-out signals on data sharing practices, specifically, the disclosure of device identifiers. A reduction in the sharing of such identifiers can indicate that an opt-out request has been respected since a third party is now more limited in associating particular data with a particular user or device.
@@
Android’s ad ecosystem is integrated into the broader online ad ecosystem and relies on the delivery of ads via the web protocol. An opt-out request, and the corresponding setting of a privacy flag, can occur in several ways. Figure 2 shows how a user’s GPC opt-out request results in setting a privacy flag that is further propagated on the web, which is the underlying protocol for most ads on mobile apps. We evaluate to which extent such flags are set via in-app opt-out interfaces, GPC signals, and disabling the AdID. It is likely that many apps do not have functionality to process GPC headers. Especially, since Android does not have a platform-level GPC setting, apps operators may not have set up server-side code to capture GPC headers. This lack of explicit support for GPC, however, does not principally change the legal status of GPC as a mandatory opt-out setting on Android and, more generally, any web-based platform.
@@ 4.3. UI Opt-out Evaluation
For determining whether the CCPA opt-out right is applicable to an app we rely on app operators’ respective declarations in their privacy policies. For our app-level UI opt-out evaluation we first chose from our app dataset 100 random apps whose operators notify California residents in their privacy policies that they have a CCPA opt-out right. To ensure that a privacy policy is applicable to an analyzed app we relied on the link from each app’s US Google Play Store Data Safety page to the policy. If necessary, we further followed links in the policy to US state- or California-specific policies. The privacy policy analysis was performed by an author with legal expertise in California privacy law and attorney admitted to the California Bar Association. Similar to previous work (Kollnig et al., 2021), we categorized each app’s opt-out setting as follows:
  • Opt-out Right Setting: App has a dedicated CCPA opt-out setting, e.g., toggle or button, to opt out of the selling or sharing of personal information per the CCPA.
  • Generic/SDK/Cookie Opt-out Setting: App has an “opt-out” setting without mentioning that it relates to the selling or sharing of personal information or that users have a right to opt out. This category covers generic opt-out settings as well as settings for cookie and SDK opt-outs.
  • Opt-out Right Form: App has a CCPA right form, e.g., that requires California residents to enter their email address, to opt out of the selling or sharing of personal information or claim other CCPA rights.
  • Privacy Policy: App only has a privacy policy link but no opt-out setting inside the app. We did not follow any privacy policy-linked opt-out setting that is only accessible via the privacy policy. If there is just a link to a “privacy policy” inside the app, it would not be clear to a user that that is where they could opt out. Such opt-out setting would not be “clear and conspicuous” and insufficient per the CCPA if the app does not respect opt-out signals via GPC.
  • No Opt-out Setting or Privacy Policy: App is missing both an opt-out setting and privacy policy link.
We evaluated whether an app has an opt-out setting, for example, under a “privacy settings” or “settings” menu. If clicking on such a menu item took us to an external website, we followed the respective procedure there. For a subset of 38 apps—those with opt-out right setting or generic/SDK/cookie opt-out setting—we performed a dynamic analysis. We did not evaluate apps with opt-out right form because submitting such form would not have immediate impact visible in the apps’ network traffic. For 31/38 apps the network captures succeeded in our two test conditions. For the remaining apps we did not collect any or only little network traffic, for example, due to apps’ crashing during the analysis.
The two test conditions under which we evaluated each app are as follows:
@@
During our evaluation we performed for each app a manual dynamic analysis. For each condition, we captured each app’s network traffic for 30 seconds, similar to previous work (Kollnig et al., 2021), using a Google Pixel 6a with Android 13, our test device. We found in preliminary tests that this time period is sufficient to understand the app’s ad tracking behavior and that the order of the conditions does not impact the results. If necessary to use the app, we created an account and logged into the app before evaluating the two test conditions. For each app, when switching conditions, we cleared the persistent storage and SharedPreferences of the test device. Upon an app’s request, we granted all permissions except for location permission as an additional measure to prevent apps from detecting our real location.
For the analysis of network traffic, we followed the state-of-the art in the literature: use of Mitmproxy (Mitmproxy Project, 2024) and circumvention of certificate pinning with Frida (Frida, 2024). For capturing an app’s network traffic, we connected the test device to a laptop using Mitmproxy in SOCKS mode (Mitmproxy Project, 2024). To issue commands to the test device, e.g., to start Mitmproxy, we connected the test device to a laptop via the Android Debug Bridge (ADB) (Google for Developers, 2024). We obtained a Californian IP address via a Mullvad VPN server (Mullvad, 2024) to ensure applicability of the CCPA. During our analysis we also manually took screenshots of the opt-out UI, if any. We performed out analysis in July and August 2025. Figure 3 shows an overview of our analysis setup.

### v6

@@ 3.2. Methods for Opting Out
The CCPA Regulations outline several methods through which businesses must allow consumers to exercise their opt-out right. [^10] CCPA Regulations §7026(a). As a general guideline, they require businesses to consider the methods by which they interact with consumers, the manner in which they collect the personal information that they make available to third parties, available technology, and ease of use by the consumer when determining which methods consumers may use to submit requests to opt out of selling and sharing with at least one method reflecting the manner in which the business primarily interacts with the consumer.
@@
For websites the CCPA requires businesses to provide a “clear and conspicuous link on the business’ internet homepages, titled ‘Do Not Sell or Share My Personal Information,’ to an internet web page that enables a consumer, or a person authorized by the consumer, to opt out of the sale or sharing of the consumer’s personal information.” [^11] CCPA §1798.135(a)(1). In this regard, cookie banners or cookie controls, by themselves, are insufficient because selling and sharing practices are not identical to cookie practices. [^12] CCPA Regulations §7026(a)(4). While there may be some overlap—for example, disabling targeting cookies may also reduce selling and sharing of personal information—the CCPA requires a dedicated method for submitting requests to opt out of the selling and sharing of personal information.
@@
In addition to links, the CCPA Regulations mandate that businesses respect opt-out preference signals: “A business that collects personal information from consumers online shall, at a minimum, allow consumers to submit requests to opt-out of sale/sharing through an opt-out preference signal and at least one of the following methods: an interactive form accessible via the ‘Do Not Sell or Share My Personal Information’ link, the Alternative Opt-out Link, or the business’s privacy policy if the business processes an opt-out preference signal in a frictionless manner.” [^13] CCPA Regulations §7026(a)(1). The “Alternative Opt-out Link” means a link combining the two separate “Do Not Sell or Share My Personal Information” and “Limit the Use of My Sensitive Personal Information” links into one “Your Privacy Choices,” or, “Your California Privacy Choices” link per CCPA Regulations §7001(b) and §7015. The idea is that businesses should (1) prominently notify consumers of the selling and sharing of personal information and how they can opt out and (2) process their opt-out preferences via automated opt-out preference signals. If a business processes opt-out preference signals in a frictionless manner, it is not required to provide any opt-out links. [^14] CCPA Regulations §7025(e). Businesses are required to treat opt-out preference signals as a valid request to opt out of the sale or sharing of personal information for a browser or device and any consumer profile associated with that browser or device, including pseudonymous profiles. [^15] CCPA Regulations §7025(c)(1).
@@ 3.3. Opting Out via GPC
The most important privacy preference signal for purposes of the CCPA is GPC, which is a draft standard at the W3C (Zimmeck et al., 2025). GPC signals are transmitted from users’ clients to servers via request headers. Servers can also query a user agent’s GPC status via the GPC DOM property. GPC can be set to 1 (header) or true (DOM property) to express a user’s opt-out as defined by applicable law in the user’s jurisdiction. Thus, GPC can be used in multiple jurisdiction and will have the meaning that a particular law or regulation attaches to it. GPC aims to enable users to exercise their opt-out rights beyond individual websites and apps for sets of sites or apps.
Recent studies evaluating the usability of GPC suggest that most users understand its purpose and that they would turn it on if it were more broadly available (Zimmeck et al., 2023, 2024b). Its rate of adoption by website operators appears to increase slowly and steadily estimated at a few percentage points per year (Hausladen et al., 2025). Compliance with GPC is required in California (since January 2021) (Becerra, 2021), Colorado (since July 2024) (Colorado Department of Law, 2024), Connecticut (Office of the Attorney General, 2025) (since January 2025), and New Jersey (New Jersey Division of Consumer Affairs, 2025) (since July 2025). Notably, in these jurisdictions the opt-out right and GPC are not limited in their applicability to websites but apply to mobile apps as well.
@@ 3.4. Opting Out via GPC on Mobile Apps
The Office of the California Attorney General brought various enforcement actions against mobile app operators for not respecting consumers’ opt-out right (State of California Department of Justice, 2022b, a, 2025). Consumers have a right to opt out on every “device,” which the CCPA defines as “any physical object that is capable of connecting to the internet, directly or indirectly, or to another device.” [^16] CCPA §1798.140(o). Thus, the International Association of Privacy Professionals, a nonprofit and non-advocacy membership association of privacy professionals, recommends for mobile apps that “[a]t a minimum, mobile apps should […] include a functional ‘Do Not Sell or Share My Personal Information’ link in the app […]” (Desai and Zweifel-Keegan, 2023). However, opting out via GPC on mobile devices presents a challenge.
The core problem is the lack of a native and platform-level GPC setting on Android. [^17] For that matter, such settings are lacking on any platform but the web. This technical gap is in conflict with the CCPA Regulations, which mandate that “[t]hrough an opt-out preference signal, a consumer can opt-out of sale and sharing of their personal information with all businesses they interact with online without having to make individualized requests with each business.” [^18] CCPA Regulations §7025(a) (emphasis added). This discrepancy between the law and the technical reality raises crucial questions: How are app operators expected to facilitate GPC opt-out signals without platform-level support. Do they even anticipate receiving such signals? On a fundamental level, as mobile apps are using the web for their client–server communication, including for ad-serving, they can generally process GPC signals in the same way as websites. Thus, our study investigates to which extent this is actually the case.
@@ 4. Methodology
Figure 2. Exercising the opt-out right via GPC. 1. An HTTP request from the user’s device is sent with a GPC header. By sending this header the user is exercising the opt-out right. 2. The HTTP response contains a script for setting a privacy flag client-side. By identifying such a privacy flag and its value we determine whether the opt-out has been respected. App operators can implement the logic for processing privacy flags themselves or use the library of a CMP. 3–4. A third-party ad network queries the privacy flag via its ad tag that the app operator integrated in its app. 5. The flag may be attached to ad requests from the user to the ad network, which is the second point that we use to determine whether the opt-out has been respected. The ad network can also act directly upon receiving GPC signals, which are attached to all requests as headers. 6. The ad network responds to ad requests from opted out users with an ad not based on sold or shared personal information.:
@@ 4.1. App Dataset
For our evaluation we downloaded a set of popular free apps from the US Google Play store in August 2023. Following previous work (Scoccia et al., 2020; Shuba and Markopoulou, 2020; Han et al., 2020), we identified the apps based on Google’s lists of popular apps in 49 app categories on the Google Play store. Each list contained 45 apps. The 49 app categories include 17 game categories (Google, 2025d). The sets of apps in the different categories are disjoint as each app can be assigned one category on the Google Play store (Google, 2025d). The 45 apps listed for each of the 49 categories are “heavily influenced by popularity” (Google, 2025c). [^19] Apps in the Communication and Productivity categories had the most installs while those in the Events category had the least. Appendix Figure 12 shows the app install ranges for the apps in all 49 app categories. We selected popular free apps for our study because those are most likely to rely on an ad-based business model that implies selling personal information or sharing such with third parties (Section 3.1). Using Raccoon (Onyxbits, 2025), we successfully downloaded 1,896/2,205 (86%) of the listed apps. The remaining apps were incompatible with our test device, removed from the Google Play store before we could download them, or not successfully downloaded. We made two download attempts per app. For our GPC opt-out evaluation we successfully performed a dynamic analysis of 1,811/1,896 (96%) apps resulting in our app dataset. For the remaining apps we did not collect any or only little network traffic, for example, due to apps’ crashing during the analysis.
@@ 4.2. Measuring Opt-out Effects
We measure the effects of opting out—i.e., whether an app stopped selling or sharing personal information—in three ways: we measure whether an app (1) sets CCPA privacy flags to opt out (such as the US Privacy String (IAB Tech Lab, 2025)), makes fewer connections to ad tracking domains, or (3) discloses fewer device identifiers to third parties. As to the first measure, it is the responsibility of ad networks to provide functionality in their ad libraries for app operators to identify ad requests to which the CCPA opt-out right applies. If an app operator enables such functionality, a privacy flag should be set on the user’s device and possibly also be attached to outgoing ad requests to third-party ad networks, as shown in Figure 2. Thus, as privacy flags are generally communicated client-to-server, for example, ad networks retrieve the US Privacy String from an Android’s device local SharedPreferences (IAB Tech Lab, 2024b), we can use client-side analysis to determine the opt-out status. If a third party receives an ad request with a privacy flag, it knows to discard personal information upon receipt and serve an ad not based on sold or shared personal information.
@@
Beyond changing privacy flag values, a successful opt-out can manifest itself through other observable changes in an app’s behavior. An app can comply by not making any ad calls. Thus, monitoring an app’s connections to third-party ad tracking domains serves as another indicator for opt-out adherence. Further, an app may continue to connect to third-party ad tracking domains but stop the sharing of identifying data. Thus, we also examine the impact of GPC signals on data sharing practices, specifically, the disclosure of device identifiers. A reduction in the sharing of such identifiers can indicate that an opt-out request has been respected since a third party is now more limited in associating particular data with a particular user or device.
@@
Android’s ad ecosystem is integrated into the broader online ad ecosystem and relies on the delivery of ads via the web protocol. An opt-out request, and the corresponding setting of a privacy flag, can be initiated in several ways. Figure 2 shows how a user initiates an opt-out request by sending a GPC header. However, it is likely that many apps do not yet have functionality to process GPC signals. Especially, since Android does not have a platform-level GPC setting, apps operators may not have set up server-side code to detect GPC signals. Thus, we also evaluate to which extent privacy flags are set via in-app opt-out interfaces or disabling the AdID. The lack of platform-level support for GPC, however, does not principally change the legal status of GPC as a mandatory opt-out setting on Android and, more generally, any web-based platform.
@@ 4.3. UI Opt-out Evaluation
In our app-level UI opt-out evaluation, for determining whether the CCPA opt-out right is applicable to an app we relied on app operators’ respective declarations in their privacy policies. We first chose from our app dataset 100 random apps whose operators notify California residents in their privacy policies that they have a CCPA opt-out right. To ensure that a privacy policy is applicable to an analyzed app we relied on the link from each app’s US Google Play store Data Safety page to the policy. If necessary, we further followed links from a policy to US state- or California-specific policies. Similar to previous work (Kollnig et al., 2021), we categorize each app’s opt-out setting as follows:
  • Opt-out Right Setting: The app has a dedicated CCPA opt-out setting, e.g., toggle or button, to opt out of the selling or sharing of personal information per the CCPA.
  • Generic/SDK/Cookie Opt-out Setting: The app has an opt-out setting without mentioning that it relates to the selling or sharing of personal information or that users have a right to opt out. This category covers generic opt-out settings as well as settings for cookie and SDK opt-outs.
  • Opt-out Right Form: The app has a CCPA right form, e.g., requiring California residents to enter their email address, to opt out of the selling or sharing of personal information or claim other CCPA rights.
  • Privacy Policy: The app only has a privacy policy link but no opt-out setting inside the app. We do not count opt-out settings that are only accessible via a privacy policy. If there is just a link to a “privacy policy” inside the app, which then links to an opt-out setting, it would not be “clear and conspicuous” and insufficient per the CCPA if the app does not respect opt-out signals via GPC.
  • No Opt-out Setting or Privacy Policy: The app is missing both an opt-out setting and a privacy policy link.
We evaluated whether an app has an opt-out setting, for example, under a “privacy settings” or “settings” menu. If clicking on such a menu item took us to an external website, we followed the respective procedure there. For a subset of 38 apps—those with opt-out right setting or generic/SDK/cookie opt-out setting—we performed a dynamic analysis. We did not evaluate apps with opt-out right form because submitting such form would not have immediate impact visible in the apps’ network traffic. For 31/38 apps we were able to capture the network traffic in our two test conditions. For the remaining apps we did not collect any or only little network traffic, for example, due to apps’ crashing during the analysis.
The two test conditions under which we evaluated each app are the following:
@@
During our evaluation we performed for each app a manual dynamic analysis. For each condition, we captured each app’s network traffic for 30 seconds, similar to previous work (Kollnig et al., 2021), using a Google Pixel 6a with Android 13, our test device. We found in preliminary tests that this time period is sufficient to understand the app’s ad tracking behavior and that the order of the conditions does not impact the results. If necessary to use the app, we created an account and logged into the app before evaluating the two test conditions. For each app, when switching conditions, we cleared the persistent storage and SharedPreferences of the test device. Upon an app’s request, we granted all permissions except for location permission as an additional measure to prevent apps from detecting our real location as we were not based in California.
For the analysis of network traffic, we followed the state-of-the art in the literature: use of Mitmproxy (Mitmproxy Project, 2025) and circumvention of certificate pinning with Frida (Frida, 2025). For capturing an app’s network traffic, we connected the test device to a laptop using Mitmproxy in SOCKS mode (Mitmproxy Project, 2025). To issue commands to the test device, e.g., to start Mitmproxy, we connected the test device to the laptop via the Android Debug Bridge (ADB) (Google for Developers, 2025). We obtained a California IP address via a Mullvad VPN server (Mullvad, 2025) to ensure applicability of the CCPA. During our analysis we also manually took screenshots of the opt-out UI, if any. We performed out analysis in July and August 2025. Figure 3 shows an overview of our analysis setup.